# Linearly-scalable and entropy-optimal learning of nonstationary and nonlinear manifolds

### v1

@@ 4 Visualising EOMC internal coordinates $\gamma$ with t-SNE
Figure 6: Nonlinear t-SNE for visualisation of the EOMC internal coordinates $\gamma$ from (2.2). Panel A: visualising $\gamma$ from the Example 1 in Sec. 3.2 (left), and $\gamma$ from the Example 2 in Sec. $3.$ Panel B: t-SNE visualisation of raw data from the L96 model described in Sec. 3.4 (left column), and for EOMC $\gamma$ (right column).
As shown above, state-of-the-art nonlinear methods like t-SNE and UMAP rely on tuning of a multitude of hyper-parameters, both method- and numerics-specific. However, in examples 1 and 2 from Sections 3.2 and 3.2, t-SNE failed to recover the low-dimensional manifolds from these nonstationary and nonlinear data (see Figs. 2B and 3B). Instead, in both of the cases it produced multiple clusters that were not present in the generated data.
Alternatively, we can deploy t-SNE to directly visualise in two or three dimensions the structures like the EOMC interpolation coefficients $\gamma$. For this, in t-SNE we can not use the common Euclidean metrics, but need to deploy a distance measure that is conform with probability measures, for example, the cross-entropy or the simmetrized Kullback-Leibler divergence. As shown in the Fig. 6A, without a particular hyper-parameter tuning (by just setting the perplexity parameter of t-SNE somewhere in the range of values between 400 and 1’200) allows fully recovering all of the original low-dimensional manifolds used in the data generation, although with some distortion. This result is a bit surprising, since it does not require taking the EOMC manifold projections from the formula (11) - and indicates that the topological structure of the manifolds is already contained in the EOMC probabilities $\gamma$.
Applying t-SNE to the raw data from the Lorenz-96 model in Sec. 3.4 reveals no pronounced clusters - although they appear to be very prominent when applying the EOMC (see Figs. 4 and 5). In contrast, without any particular hyper-parameter tuning, t-SNE with the symmetrized KL-distance reveals a very pronounced cluster structure in the $\gamma$ from EOMC. In accordance with the other results obtained in the Sec. 3.4, t-SNE reveals that the number of low-dimensional manifold clusters gradually grows with the growth of the L96 extern forcing $F$ - but the topology of these manifolds looks very similar, differing only in their position and orientation.
@@ 5 Discussion
Data-driven manifold learning and dimensionality reduction methods are constituting one of the central pillars of data analysis in many areas of science. For example, in neurosciences and bioinformatics, nonlinear methods like t-SNE and UMAP are very widely used to investigate and to visualise the high-dimensional data structures, as well as to detect clusters. However, examples provided above illustrate that t-SNE - one of the most popular tools, with over 28’000 citations according to Google Scholar - can create clusters that are not present in the data (see Figs. 2B and 3B), as well as to not detect clusters when they are actually present (see the left panels of Fig. 6B). Another problems of such methods include polynomial, or, in the best cases, $\mathcal{O}\left(T\log\left(T\right)\right)$ cost scaling with the size $T$ of data statistics - as well as necessity to tune multiple model-specific and numerics-specific hyper-parameters, and with no consensus on the optimal procedure needed to select them..
As demonstrated in the Sec. 2.2 and proven in the Lemmas 3, 4, and in the Theorem 1, Entropy Optimal Manifold Clustering (EOMC) allows mitigating these problems, resulting in computational cost scaling of $\mathcal{O}\left(T\right)$, with an explicit rule (11) to project the new data points on the manifold with cost scaling of $\mathcal{O}\left(D\right)$ in the leading order, providing a very robust learning of very nonlinear manifolds from noisy and nonstationary data, and not requiring complicated hyper-parameter adjustment (as shown on the synthetic examples the Sec. 3.2 and 3.3).
@@
However, the results obtained in the Sec. 3.4 tell a somewhat different story. EOMC analysis reveals that the internal coordinates $\gamma$ in all of the considered forcing regimes, exhibit very persistent and metastable dynamics in time, indicating relatively long stays in each of the local linear manifolds (with relatively low dimensionality $d=3$). As was analysed above, this metastability of $\gamma$ can not be an artefact of the EOMC data analysis: as can be seen from (2.2), the EOMC loss function does not contain any terms that would enforce the persistence or metastability on $\gamma$. As a matter of fact, the value of $\mathcal{L}^{\textrm{EOMC}}$ in (2.2) is invariant with respect to any permutation of the columns of data matrix $X$. This means that the observed metastability of $\gamma$ for Lorenz-96 in chaotic and very chaotic regimes can only be an imprint of the underlying L96 dynamics - that appears to be best described by a metastable process switching between low-dimensional (with $d=3$) locally-linear manifolds. Analysis from Sec. 3.4 revealed that the main effect of increasing forcing was in gradually-increasing the total number $K$ of these local low-dimensional manifolds - and in slowly decreasing mean relaxation and mean exit times, that were measured from the transfer operators inferred from EOMC variables $\gamma$ (see Fig. 5). We shown that average Markovian mean exit times $\mathbb{E}_{k}\left[\tau_{k}^{ex}(P)\right]$ closely match the doubling times from positive Lyapunov exponents in the literature (see Fig. 5B) - indicating the predictability horizons quantified with Lyapunov exponents as limits for the average times it takes for the system to leave the current low-dimensional manifold and to go somewhere else. And, it was shown that the transfer operator description revealed by the EOMC analysis in Figs. 5 and 6 allows going beyond this limit: besides allowing to measure the average time $\tau_{k}^{ex}(P)$ the dynamics spends in the local manifold $k$, transfer operator provides the transition probabilities $P_{j,k}$ to all of the other local manifold states $j\neq k$, where the dynamics can go to after leaving $k$. For Markov processes, relaxation times define the predictability horizon of the system, quantifying the time that the system requires to forget its initial condition [34]. As revealed by the relaxation times curve $\tau^{rel}(P)$ (see the blue dotted curve in the Fig. 5B), this transfer operator description roughly doubles the prediction horizon for the system, as compared to the state of the art Lyapunov exponent description.

### v2

@@ 3.2 Analysis of data from Lorenz-63 in chaotic, strongly-chaotic and very strongly-chaotic regimes
3.3 Analysis of data from modified Hasegawa-Wakatani (mHW) model of drift-wave turbulence in the edge of a tokamak plasma.
One-dimensional L96-model is considered by many to provide only a very limited, and simplified description of turbulent atmospheric dynamics - and not a model of a "real" turbulence behaviour. To check if the findings from previous Section about metastability and predictability horizons of L96 are induced by this oversimplification - or if they are also reflecting intrinsic properties of "real" turbulent systems, next we will consider an application of EOMC to the output of a more realistic model from Magnetohydrodynamics (MHD). We will take the Hasegawa-Wakatani model - a seminal two-field fluid description model of drift-wave turbulence in the edge of a tokamak plasma [38, 39, 40]. It reduces the complex MHD equations to the evolution of density fluctuations ($n$) and electrostatic potential ($\phi$) fields.ăThe modified Hasegawa-Wakatani (mHW) model describes the evolution of electrostatic potential $\phi$ and density fluctuations $n$ in a two-dimensional slab geometry. In the following application example, we will deploy the mHW model version introduced by Numata et al. [41], involving the resistive coupling term acting only on non-zonal fluctuations.
Governing Equations
ă Let the zonal average of a field $f$ be defined as $\langle f\rangle=\frac{1}{L_{y}}\int f\,dy$, and the non-zonal fluctuation as $\tilde{f}=f-\langle f\rangle$. The mHW equations are:
$\frac{\partial\zeta}{\partial t}+\{\phi,\zeta\}$ $=\alpha(\tilde{\phi}-\tilde{n})-\mu\nabla^{4}\zeta$ (17)
$\frac{\partial n}{\partial t}+\{\phi,n\}+\kappa\frac{\partial\phi}{\partial y}$ $=\alpha(\tilde{\phi}-\tilde{n})-\mu\nabla^{4}n$ (18)
where $\zeta=\nabla^{2}\phi$ is the ion vorticity and $\{a,b\}=\partial_{x}a\partial_{y}b-\partial_{y}a\partial_{x}b$ is the Poisson bracket representing $\mathbf{E}\times\mathbf{B}$ advection. Variables and parameters of mHW model equations [41] have the following meaning:
  • $\phi$: electrostatic potential;
  • $n$: electron density fluctuations;
  • $\zeta$: ion vorticity ($\nabla^{2}\phi$);
  • $\alpha$: adiabaticity parameter (resistive coupling);
  • $\kappa$: background density gradient scale length ($\partial_{x}\ln n_{0}$);
  • $\mu$: dissipation/viscosity coefficient.
Data generation
To generate the time series for the EOMC analysis, we use the MATLAB code by Jean-Christophe Nave and Denis St-Onge, available at https://github.com/DenSto/HWE_solver. We use the same settings for all of the mWH model parameters as in this code, with the only change being a slightly reduced grid size ($64\times 64$ grid points). We generate a time series on the interval $\left[600,3000\right]$ (skipping the outputs between $t=0$ and $t=600$, when the model "burns-in"), with constant time intervals $\delta t=0.3125$. Finally, we create a data matrix $X$ for EOMC, containing the time series instances for electron density fluctuation $n$ and ion vorticity $\zeta$ fields (see Fig. 7A).
Figure 7: Application of EOMC to the time series output of modified 2D Hasegawa-Wakatani model (mHW) [38, 39, 40]. For details of the description see Sec. $3.$
Application of EOMC to mHW output data
We follow the same analysis procedure as deployed for the previous L96-example in Sec. 3.2: first, we determine the optimal values for the reduced manifold dimension $d$ (that appears to be $d=4$), and $K$ (that appears to be $K=80$, see Fig. 7B). The optimal values for the other hyper-parameters appear to be $\alpha=0.1,\beta=10^{-3},\beta_{0}=1.0$ - and as in the previous examples, obtained results appear to remain robust in broad ranges of these hyper-parameter values. As can be seen when comparing Fig. 7B with the middle column panels of Fig. 5, similarly to L96, for the same level of lossy data compression, application of EOMC to mHW data allows achieving over around 15-fold smaller relative lossy compression errors then PCA. Fig. 7C shows that the optimal $\gamma(k,t)$ appears to be a discrete, metastable and persistent transition process - staying long times in the same local low-dimensional manifolds before switching to the next ones. Finally - and very similarly to the previous L96-example (see Fig. 5 and 6) - Markovian transition probability matrix $P$ obtained from $\gamma(k,t)$ appears to be very metastable, with mean exit times of 5.9 and the predictability horizon (measured as the mean relaxation time of $P$) being almost twice as long - around 10.8 time units.
@@ 4 Visualising EOMC internal coordinates $\gamma$ with t-SNE
Figure 8: Nonlinear t-SNE for visualisation of the EOMC internal coordinates $\gamma$ from (2.2). Panel A: visualising $\gamma$ from the Example 1 in Sec. 3.1.1 (left), and $\gamma$ from the Example 2 in Sec. 3.1.2. Panel B: t-SNE visualisation of raw data from the L96 model described in Sec. 3.2 (left column), and for EOMC $\gamma$ (right column).
As shown above, state-of-the-art nonlinear methods like t-SNE and UMAP rely on tuning of a multitude of hyper-parameters, both method- and numerics-specific. However, in examples 1 and 2 from Sections 3.1.1 and 3.1.1, t-SNE failed to recover the low-dimensional manifolds from these nonstationary and nonlinear data (see Figs. 2B and 3B). Instead, in both of the cases it produced multiple clusters that were not present in the generated data.
Alternatively, we can deploy t-SNE to directly visualise in two or three dimensions the structures like the EOMC interpolation coefficients $\gamma$. For this, in t-SNE we can not use the common Euclidean metrics, but need to deploy a distance measure that is conform with probability measures, for example, the cross-entropy or the symmetrized Kullback-Leibler divergence. As shown in the Fig. 8A, without a particular hyper-parameter tuning (by just setting the perplexity parameter of t-SNE somewhere in the range of values between 400 and 1’200) allows fully recovering all of the original low-dimensional manifolds used in the data generation, although with some distortion. This result is a bit surprising, since it does not require taking the EOMC manifold projections from the formula (11) - and indicates that the topological structure of the manifolds is already contained in the EOMC probabilities $\gamma$.
Applying t-SNE to the raw data from the Lorenz-96 model in Sec. 3.2 reveals no pronounced clusters - although they appear to be very prominent when applying the EOMC (see Figs. 5 and 6). In contrast, without any particular hyper-parameter tuning, t-SNE with the symmetrized KL-distance reveals a very pronounced cluster structure in the $\gamma$ from EOMC. In accordance with the other results obtained in the Sec. 3.2, t-SNE reveals that the number of low-dimensional manifold clusters gradually grows with the growth of the L96 extern forcing $F$ - but the topology of these manifolds looks very similar, differing only in their position and orientation.
Applying t-SNE with a symmetrized Kullback-Leibler divergence as distance to the EOMC interpolation coefficients $\gamma$ from mHW example analysis in Sec. 3.3 results in a very similar picture - with multiple clearly distinct clusters (see Fig. 7E) of very similar topology, differing only in their position and orientation.
@@ 5 Discussion
Data-driven manifold learning and dimensionality reduction methods are constituting one of the central pillars of data analysis in many areas of science. For example, in neurosciences and bioinformatics, nonlinear methods like t-SNE and UMAP are very widely used to investigate and to visualise the high-dimensional data structures, as well as to detect clusters. However, examples provided above illustrate that t-SNE - one of the most popular tools, with over 28’000 citations according to Google Scholar - can create clusters that are not present in the data (see Figs. 2B and 3B), as well as to not detect clusters when they are actually present (see the left panels of Fig. 8B). Another problems of such methods include polynomial, or, in the best cases, $\mathcal{O}\left(T\log\left(T\right)\right)$ cost scaling with the size $T$ of data statistics - as well as necessity to tune multiple model-specific and numerics-specific hyper-parameters, and with no consensus on the optimal procedure needed to select them. Last but not least, state-of-the-art manifold learning and nonlinear dimensionality reduction methods do not provide straightforwardly-computable measures of input data reliability - reducing robustness of these approaches and making them vulnerable to the adversarial attacks.
As demonstrated in the Sec. 2.2 and proven in the Lemmas 3, 4, and in the Theorem 1, Entropy Optimal Manifold Clustering (EOMC) allows mitigating these problems, resulting in computational cost scaling of $\mathcal{O}\left(T\right)$, with an explicit rule (11) to project the new data points on the manifold with cost scaling of $\mathcal{O}\left(D\right)$ in the leading order, providing a very robust learning of very nonlinear manifolds from noisy and nonstationary data, not requiring complicated hyper-parameter adjustment (as shown on the synthetic examples the Sec. 3.1.1 and 3.1.2). Moreover, as was shown in the Sec. 2.3, deploying the explicit analytical solution (15), one can easily modify EOMC without an increase of leading order computational cost - and allow simultaneous learning of the data reliability measure $\gamma_{0}$. As demonstrated in Sec. 3.1.3 and Fig. 4, this modification of EOMC allows a much more robust and artefact-free functioning of the method (compare Fig. 4B to Fig. 4C).
@@
From the current perspective, predictability in weather systems or inhomogeneous turbulence / geophysical fluids more generally, is typically limited by the emergence of coherent structures whose growth rates are local (in space and time). Predictability in the observed weather systems is then considered to be a local property, with all the complexity that it brings. For the high-dimensional dynamical systems, Lyapunov exponents and local Kaplan-Yorke dimensions are considered to be the best currently-available measures of local dimensionality and predictability [43].
The results obtained in the Sec. 3.2 tell a somewhat different story. EOMC analysis reveals that the internal coordinates $\gamma$ in all of the considered forcing regimes, exhibit very persistent and metastable dynamics in time, indicating relatively long stays in each of the local linear manifolds (with relatively low dimensionality $d=3$). Similar metastable behaviour for a simplified barotropic quasi-geostrophic ODE model of atmospheric dynamics was previously demonstrated applying Hidden Markov Model (HMM) analysis from machine learning, and reported in the work of A. Majda et al [44]. However, HMMs impose a Markovian structure and Markov-property on the data during analysis procedure (i.e., the property that every next observation in time series is conditionally-dependent on the previous one). The value of Markovian functional used in the HMM training is not invariant with respect to permutations of the data, so also the order of the data in time series matters for the HMM analysis outcome. In another words, the metastability of a simplified barotropic quasi-geostrophic ODE model reported in [44] could be alternatively explained as an artefact of HMM data analysis.
@@
In contrast, as was analysed above in Sec. 3.2, metastability of $\gamma$ can not be an artefact of the EOMC data analysis: as can be seen from (2.2), the EOMC loss function does not contain any terms that would enforce the persistence, metastability or Markov-property on $\gamma$. As a matter of fact, the value of $\mathcal{L}^{\textrm{EOMC}}$ in (2.2) is invariant with respect to any permutation of the columns of data matrix $X$. This means that the observed metastability of $\gamma$ for Lorenz-96 in chaotic and very chaotic regimes can only be an imprint of the underlying L96 dynamics - that appears to be best described by a metastable process switching between low-dimensional (with $d=3$) locally-linear manifolds. Analysis from Sec. 3.2 revealed that the main effect of increasing forcing was in gradually-increasing the total number $K$ of these local low-dimensional manifolds - and in slowly decreasing mean relaxation and mean exit times, that were measured from the transfer operators inferred from EOMC variables $\gamma$ (see Fig. 6). We shown that average Markovian mean exit times $\mathbb{E}_{k}\left[\tau_{k}^{ex}(P)\right]$ closely match the doubling times from positive Lyapunov exponents in the literature (see Fig. 6B) - indicating the predictability horizons quantified with Lyapunov exponents as limits for the average times it takes for the system to leave the current low-dimensional manifold and to go somewhere else. And, it was shown that the transfer operator description revealed by the EOMC analysis in Figs. 5 and 6 allows going beyond this limit: besides allowing to measure the average time $\tau_{k}^{ex}(P)$ the dynamics spends in the local manifold $k$, transfer operator provides the transition probabilities $P_{j,k}$ to all of the other local manifold states $j\neq k$, where the dynamics can go to after leaving $k$. For Markov processes, relaxation times define the predictability horizon of the system, quantifying the time that the system requires to forget its initial condition [34]. As revealed by the relaxation times curve $\tau^{rel}(P)$ (see the blue dotted curve in the Fig. 6B), this transfer operator description roughly doubles the prediction horizon for the system, as compared to the state of the art Lyapunov exponent description.
However, one-dimensional L96-model provides only a very limited, and a very simplified description of turbulent atmospheric dynamics. To check if the findings about metastability and predictability horizons of L96 are induced solely by this oversimplification - or if they are also reflecting intrinsic properties of "real" turbulent systems, next we considered an application of EOMC to the output of a more realistic model from Magnetohydrodynamics (MHD): we took the modified Hasegawa-Wakatani model (mHW) - a seminal two-field fluid description model of drift-wave turbulence in the edge of a tokamak plasma [38, 39, 40]. As can be seen from the results provided in the Sec. 3.3 and from comparisons of Figs. 5 and 6 to a Fig. 7, results for this much more advanced mHW model tell basically the same story as the L96 results: EOMC reveals that the dynamics is best represented by a metastable process infrequently switching between the different local low-dimensional manifolds, with the predictability horizon given by the mean relaxation times, almost doubling the prediction horizons measured with the mean exit times and Lyapunov exponents. In another words, metastability of topology and dynamics is not an artefact of L96-simplification, but is an essential characteristics of much more complex turbulent models like mHW, that do not rely on L96-simplifications.